### v1

@@ 3.2 Distances in $G$ and $H$
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[k-1,j]$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(G)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma 5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Hence, $d_{G}\left(x_{k},x_{j}\right)\leq k-j+4$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm 1 of Algorithm 1, there exist $u\in S_{j+1}$ and $w,v\in S_{j}$ such that $\{x^{\prime}_{j}w,wu,uv,vx_{j}\}\subseteq E(G)$. Hence, $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 4$ and thus, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+4$.
@@
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition 1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Algorithm 1 ensures that $x_{i}\in C^{1}_{i}$. Observe that $x^{\prime}_{i+1}\in D^{1}_{i}$. Moreover, any vertex in $S_{i}$ which is adjacent to $x^{\prime}_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x^{\prime}_{i}\in C^{1}_{i}$. Since $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $G$, and $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$. Due to 13.1, length of $R_{1}$ is at most $k-j+4$. Length of $R_{2}$ is $j-\ell$. Due to 13.2, $x_{\ell}x^{\prime}_{\ell}\in E(G)$. Hence $d_{G}\left(x,y\right)\leq d_{H}\left(x,y\right)+5$. ∎
@@ 3.3 Induced cycles in $H$
The main purpose of this section is to prove that any cluster of $G$ contains at most two vertices of an induced cycle in $H$ (see Lemma 16). We start by proving the following.
@@
Assume there exists an integer $k$ and two vertices $u_{1},u_{2}\in V(C)\cap N_{k}(r)$ such that $u_{1},u_{2}$ lie in different clusters $S_{1},S_{2}$ respectively. Let $S$ be the least common ancestor of $S_{1}$ and $S_{2}$ in $\Gamma\left({r},{G}\right)$. Observe that $S,S_{1},S_{2}$ are all distinct clusters. For $i\in\{1,2\}$, let $S^{\prime}_{i}$ denote the cluster which is a child of $S$ and lie in the $(S,S_{i})$-path in $\Gamma\left({r},{G}\right)$. Observe that $S,S^{\prime}_{1},S^{\prime}_{2}$ are distinct clusters. Since there are two $(u_{1},u_{2})$-paths in $C$ (and thus in $H$) which are internally vertex disjoint, due to 2, $S$ must contain two vertices $v_{1},v_{2}\in V(C)$. Observe that for $i\in\{1,2\},j\in\{1,2\}$, $v_{i}$ must have a neighbor in $C$ (and thus in $H$) that lies in $S^{\prime}_{j}$. Now applying 3(b) on $S^{\prime}_{1}$, we have that $G\left[P\left(r,S^{\prime}_{1}\right)\right]$ has two connected components $C_{1},D_{1}$. Similarly, applying 3(b) on $S^{\prime}_{2}$, we have that $G\left[P\left(r,S^{\prime}_{2}\right)\right]$ has two connected components $C_{2},D_{2}$. Without loss of generality, $v_{1}\in C_{1}\cap C_{2}$ and $v_{2}\in D_{1}\cap D_{2}$. But this contradicts Lemma 6. ∎
@@
Due to Lemma 8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Let $S_{k}\subseteq N_{k}(r)$ and $S_{\ell}\subseteq N_{\ell}(r)$ denote the clusters that contain some vertices of $C$. Due to Lemma 14, $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{\ell}(r)\subseteq S_{\ell}$. Now, $(a)$ follows from 4.
To prove $(b)$, assume $\{a,b,c\}\subseteq S_{\ell}\cap V(C)$ and $F=\{uv\in E(H)\colon\{u,v\}\subseteq\{a,b,c\}\}$. By 5(a), $|F|\leq 1$. Due to Lemma 14, there exists a child $S^{\prime}$ of $S_{\ell}$ in $\Gamma\left({r},{G}\right)$ such that for each $d\in\{a,b,c\}$ there exists $d^{\prime}\in S^{\prime}$ such that $dd^{\prime}\in E(H)$. This contradicts 3(a). ∎
@@
(a) Notations used in the proof of 16.1.
@@
(b) Notations used in the proof of 16.2.
@@
(c) Notations used in the proof of 16.3.
Figure 3: Illustration of the proof of Lemma 16. The dashed lines indicate edges in $H$. The solid lines indicate edges in $G$. The curved lines indicate induced paths in $G$. The vertices in the shaded region indicate cliques in $G$. There are no edges in $G$ between a vertex in a rectangular region and a vertex in a elliptical region.
@@
Let $k=\max\{d_{H}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{H}\left(r,V(C)\right)$. Due to Lemma 8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Due to Lemma 14, there are clusters $S_{k}$ and $S_{\ell}$ such that $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{k}(r)\subseteq S_{\ell}$. Due to Lemma 15, we have that $|S_{k}\cap V(C)|\leq 2$ and $|S_{\ell}\cap V(C)|\leq 2$. Moreover, all vertices of $C$ lie in the clusters that lie in the unique $(S_{k},S_{\ell})$-path in $\Gamma\left({r},{G}\right)$.
Assume the lemma is not true, and let $k^{\prime}$ be the maximum integer such that there exists a cluster $S\subseteq N_{r}(k^{\prime})$ with $|V(C)\cap S|\geq 3$. Due to Lemma 15, $\ell<k^{\prime}<k$. Let $T$ be the child of $S$ that lie in the unique $(S_{k},S)$-path in $\Gamma\left({r},{G}\right)$, and $U$ be the parent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $U\subseteq N_{k-1}(r)$ and $T\subseteq N_{k+1}(r)$. Due to 3(a), there exists at least one vertex $a\in V(C)\cap S$ such that $T$ does not contain any vertex of $C$ which is adjacent to $a$. Let $a_{1},a_{2}$ be the two neighbors of $a$ in $C$. We prove the following claim.
@@
Assuming otherwise, let $\{a_{1},a_{2}\}\subseteq U$. See Figure 3(a) for notations. Since $U$ is the parent of $S$ in $\Gamma\left({r},{G}\right)$, $P\left(r,S\right)\subseteq U$. Now, $P\left(r,S\right)$ contains two vertices $a_{1},a_{2}$ which is adjacent to some vertices (e.g. $a$) in $S$. Due to 3(b), we have that $G\left[P\left(r,S\right)\right]$ have two connected components $C_{1},C_{2}$ such that $a_{1}\in C_{1},a_{2}\in C_{2}$. Observe that, $C_{1},C_{2}$ induce cliques in $G$ and that no vertex of $U\setminus\{a_{1},a_{2}\}$ is adjacent to any vertex of $S$.
@@
Now apply 1 on each of the edges in $\{aa_{1},aa_{2},b_{1}a_{1},b_{2}a_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},a_{2}^{\prime},a^{\prime\prime}_{1},a^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},aa^{\prime}_{2},b_{1}a^{\prime\prime}_{1},b_{2}a^{\prime\prime}_{2}\}\subseteq E(G)$ and $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq C_{1}$ and $\{a^{\prime}_{2},a^{\prime\prime}_{2},a_{2}\}\subseteq C_{2}$. Similarly, apply 1 on each of the edges in $\{b_{1}c_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},b^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c_{1},b^{\prime}_{2}c_{2},b_{1}b^{\prime}_{1},b_{2}b^{\prime}_{2}\}\subseteq E(G)$.
Since $c_{1},c_{2}$ lies in the same cluster $T\subseteq N_{k+1}(r)$, there exists a $(c_{1},c_{2})$-induced path $Q$ in $G-B_{k}\left({r}\right)$. On the other hand there exists a $(a_{1},a_{2})$-upper path (see Section 2.2 for definition) $R$ in $G$ such that $(V(R)\setminus\{a_{1},a_{2}\})\subseteq B_{k-2}\left({r}\right)$. Now consider the vertices
@@
It can be verified that that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition 2. ∎
Due to 16.1, we assume (without loss of generality) $a_{1}\in U,a_{2}\in S$. For $i\in\{1,2\}$, let $b_{i}$ be the neighbor of $a_{i}$ in $C$ which is distinct from $a$. Now we prove the following claim.
@@
Assume $b_{2}\in U$. See Figure 3(b) for notations. Due to 3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,b_{2}\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},b_{2}\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exist a $(x,a_{1})$-path $P_{1}\subseteq C$ that does not contain $a$ or $a_{2}$. Similarly, there exists a $(x,b_{2})$-path $P_{2}\subseteq C$ that does not contain $a,a_{1}$ or $a_{2}$. Moreovoer, $V(P_{1})\cap V(P_{2})=\{x\}$ and the edge $a_{1}b_{1}$ lies in $E(P_{1})$. Let $c_{2}$ be the neighbor of $b_{2}$ in $P_{2}$. Due to 3, we have that $b_{1}\in S$ and $c_{2}\in S$.
Let $d_{1}\in V(P_{1})$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Let $d_{2}\in V(P_{2})$ be the neighbor of $c_{2}$ distinct from $b_{2}$. For each $i\in\{1,2\}$, observe that $d_{i}\in T$, otherwise $d_{i}$ will be adjacent to $a_{1}$ or $b_{2}$ contradicting the fact that $C$ is an induced cycle. Now apply 1 on each of the edges in $\{aa_{1},a_{2}b_{2},b_{1}a_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},b_{2}^{\prime},a^{\prime\prime}_{1},b^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}a^{\prime}_{2},b_{1}a^{\prime\prime}_{1},c_{2}b^{\prime\prime}_{2}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{b^{\prime}_{2},b^{\prime\prime}_{2},b_{2}\}\subseteq B$. Similarly, apply 1 on edges in $\{b_{1}d_{1},c_{2}d_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},c^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}d_{1},d_{2}c^{\prime}_{2},b_{1},b^{\prime}_{1},c_{2}c^{\prime}_{2}\}\subseteq E(G)$.
@@
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition 2. ∎
@@
Assume $b_{2}\in T$. See Figure 3(c) for notations. Let $z$ be the vertex in $U$ which is also adjacent to $a_{2}$ in $H$. Due to 5, we have that $z$ is distinct from $a_{1}$. Moreover, due to 3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,z\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},z\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exists a $(x,a_{1})$-path $P\subseteq C$ that does not contain $a$ or $a_{2}$. Moreover, $a_{1}b_{1}\in E(P)$ and $z\notin V(P)$. Due to 3, we have that $b_{1}\in S$. Let $c\in V(P)$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Since $H$ is triangle-free, $c$ is not adjacent to $a_{1}$. Due to 5, we have $c\neq z$ and moreover, $c$ is not adjacent to $z$ in $H$. Hence, $c\notin U,c\notin S$. Therefore, $c$ must lie in $T$.
Now apply 1 on each of the edges in $\{aa_{1},a_{2}z,b_{1}a_{1}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},z^{\prime},a^{\prime\prime}_{1}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}z^{\prime},b_{1}a^{\prime\prime}_{1}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{z,z^{\prime}\}\subseteq B$. Similarly, apply 1 on edges in $\{b_{1}c,a_{2}b_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},a^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c,a^{\prime}_{2}b_{2},b_{1}b^{\prime}_{1},a_{2}a^{\prime}_{2}\}\subseteq E(G)$.
@@
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition 2. ∎
Claims 16.1, 16.2, and 16.3 imply that $\{a,a_{2},b_{2}\}\subseteq S$. Since $\{aa_{2},a_{2}b_{2}\}\in E(H)$, we have a contradiction due to 5. ∎

### v2

@@ 3.2 Distances in $G$ and $H$
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[k-1,j]$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(G)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma˜5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Hence, $d_{G}\left(x_{k},x_{j}\right)\leq k-j+4$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm˜1 of Algorithm˜1, there exist $u\in S_{j+1}$ and $w,v\in S_{j}$ such that $\{x^{\prime}_{j}w,wu,uv,vx_{j}\}\subseteq E(G)$. Hence, $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 4$ and thus, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+4$.
@@
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition˜1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Algorithm˜1 ensures that $x_{i}\in C^{1}_{i}$. Observe that $x^{\prime}_{i+1}\in D^{1}_{i}$. Moreover, any vertex in $S_{i}$ which is adjacent to $x^{\prime}_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x^{\prime}_{i}\in C^{1}_{i}$. Since $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $G$, and $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$. Due to ˜13.1, length of $R_{1}$ is at most kj+4k-j+4. Length of $R_{2}$ is $j$-$\ell$. Due to ˜13.2, $x_{\ell}x^{\prime}_{\ell}\in E(G)$. Hence $d_{G}\left(x,y\right)\leq d_{H}\left(x,y\right)+5$. ∎
@@ 3.3 Induced cycles in $H$
The main purpose of this section is to prove that any cluster of $G$ contains at most two vertices of an induced cycle in $H$ (see Lemma˜16). We start by proving the following.
@@
Assume there exists an integer $k$ and two vertices $u_{1},u_{2}\in V(C)\cap N_{k}(r)$ such that $u_{1},u_{2}$ lie in different clusters $S_{1},S_{2}$ respectively. Let $S$ be the least common ancestor of $S_{1}$ and $S_{2}$ in $\Gamma\left({r},{G}\right)$. Observe that $S,S_{1},S_{2}$ are all distinct clusters. For $i\in\{1,2\}$, let $S^{\prime}_{i}$ denote the cluster which is a child of $S$ and lie in the $(S,S_{i})$-path in $\Gamma\left({r},{G}\right)$. Observe that $S,S^{\prime}_{1},S^{\prime}_{2}$ are distinct clusters. Since there are two $(u_{1},u_{2})$-paths in $C$ (and thus in $H$) which are internally vertex disjoint, due to ˜2, $S$ must contain two vertices $v_{1},v_{2}\in V(C)$. Observe that for $i\in\{1,2\},j\in\{1,2\}$, $v_{i}$ must have a neighbor in $C$ (and thus in $H$) that lies in $S^{\prime}_{j}$. Now applying ˜3(b) on $S^{\prime}_{1}$, we have that $G\left[P\left(r,S^{\prime}_{1}\right)\right]$ has two connected components $C_{1},D_{1}$. Similarly, applying ˜3(b) on $S^{\prime}_{2}$, we have that $G\left[P\left(r,S^{\prime}_{2}\right)\right]$ has two connected components $C_{2},D_{2}$. Without loss of generality, $v_{1}\in C_{1}\cap C_{2}$ and $v_{2}\in D_{1}\cap D_{2}$. But this contradicts Lemma˜6. ∎
@@
Due to Lemma˜8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Let $S_{k}\subseteq N_{k}(r)$ and $S_{\ell}\subseteq N_{\ell}(r)$ denote the clusters that contain some vertices of $C$. Due to Lemma˜14, $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{\ell}(r)\subseteq S_{\ell}$. Now, $(a)$ follows from ˜4.
To prove $(b)$, assume $\{a,b,c\}\subseteq S_{\ell}\cap V(C)$ and $F=\{uv\in E(H)\colon\{u,v\}\subseteq\{a,b,c\}\}$. By ˜5(a), $|F|\leq 1$. Due to Lemma˜14, there exists a child $S^{\prime}$ of $S_{\ell}$ in $\Gamma\left({r},{G}\right)$ such that for each $d\in\{a,b,c\}$ there exists $d^{\prime}\in S^{\prime}$ such that $dd^{\prime}\in E(H)$. This contradicts ˜3(a). ∎
@@
(a) Notations used in the proof of ˜16.1.
@@
(b) Notations used in the proof of ˜16.2.
@@
(c) Notations used in the proof of ˜16.3.
Figure 3: Illustration of the proof of Lemma˜16. The dashed lines indicate edges in $H$. The solid lines indicate edges in $G$. The curved lines indicate induced paths in $G$. The vertices in the shaded region indicate cliques in $G$. There are no edges in $G$ between a vertex in a rectangular region and a vertex in a elliptical region.
@@
Let $k=\max\{d_{H}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{H}\left(r,V(C)\right)$. Due to Lemma˜8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Due to Lemma˜14, there are clusters $S_{k}$ and $S_{\ell}$ such that $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{k}(r)\subseteq S_{\ell}$. Due to Lemma˜15, we have that $|S_{k}\cap V(C)|\leq 2$ and $|S_{\ell}\cap V(C)|\leq 2$. Moreover, all vertices of $C$ lie in the clusters that lie in the unique $(S_{k},S_{\ell})$-path in $\Gamma\left({r},{G}\right)$.
Assume the lemma is not true, and let $k^{\prime}$ be the maximum integer such that there exists a cluster $S\subseteq N_{r}(k^{\prime})$ with $|V(C)\cap S|\geq 3$. Due to Lemma˜15, $\ell<k^{\prime}<k$. Let $T$ be the child of $S$ that lie in the unique $(S_{k},S)$-path in $\Gamma\left({r},{G}\right)$, and $U$ be the parent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $U\subseteq N_{k-1}(r)$ and $T\subseteq N_{k+1}(r)$. Due to ˜3(a), there exists at least one vertex $a\in V(C)\cap S$ such that $T$ does not contain any vertex of $C$ which is adjacent to $a$. Let $a_{1},a_{2}$ be the two neighbors of $a$ in $C$. We prove the following claim.
@@
Assuming otherwise, let $\{a_{1},a_{2}\}\subseteq U$. See Figure˜3(a) for notations. Since $U$ is the parent of $S$ in $\Gamma\left({r},{G}\right)$, $P\left(r,S\right)\subseteq U$. Now, $P\left(r,S\right)$ contains two vertices $a_{1},a_{2}$ which is adjacent to some vertices (e.g. $a$) in $S$. Due to ˜3(b), we have that $G\left[P\left(r,S\right)\right]$ have two connected components $C_{1},C_{2}$ such that $a_{1}\in C_{1},a_{2}\in C_{2}$. Observe that, $C_{1},C_{2}$ induce cliques in $G$ and that no vertex of $U\setminus\{a_{1},a_{2}\}$ is adjacent to any vertex of $S$.
@@
Now apply ˜1 on each of the edges in $\{aa_{1},aa_{2},b_{1}a_{1},b_{2}a_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},a_{2}^{\prime},a^{\prime\prime}_{1},a^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},aa^{\prime}_{2},b_{1}a^{\prime\prime}_{1},b_{2}a^{\prime\prime}_{2}\}\subseteq E(G)$ and $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq C_{1}$ and $\{a^{\prime}_{2},a^{\prime\prime}_{2},a_{2}\}\subseteq C_{2}$. Similarly, apply ˜1 on each of the edges in $\{b_{1}c_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},b^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c_{1},b^{\prime}_{2}c_{2},b_{1}b^{\prime}_{1},b_{2}b^{\prime}_{2}\}\subseteq E(G)$.
Since $c_{1},c_{2}$ lies in the same cluster $T\subseteq N_{k+1}(r)$, there exists a $(c_{1},c_{2})$-induced path $Q$ in $G-B_{k}\left({r}\right)$. On the other hand there exists a $(a_{1},a_{2})$-upper path (see Section˜2.2 for definition) $R$ in $G$ such that $(V(R)\setminus\{a_{1},a_{2}\})\subseteq B_{k-2}\left({r}\right)$. Now consider the vertices
@@
It can be verified that that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
Due to ˜16.1, we assume (without loss of generality) $a_{1}\in U,a_{2}\in S$. For $i\in\{1,2\}$, let $b_{i}$ be the neighbor of $a_{i}$ in $C$ which is distinct from $a$. Now we prove the following claim.
@@
Assume $b_{2}\in U$. See Figure˜3(b) for notations. Due to ˜3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,b_{2}\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},b_{2}\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exist a $(x,a_{1})$-path $P_{1}\subseteq C$ that does not contain $a$ or $a_{2}$. Similarly, there exists a $(x,b_{2})$-path $P_{2}\subseteq C$ that does not contain $a,a_{1}$ or $a_{2}$. Moreovoer, $V(P_{1})\cap V(P_{2})=\{x\}$ and the edge $a_{1}b_{1}$ lies in $E(P_{1})$. Let $c_{2}$ be the neighbor of $b_{2}$ in $P_{2}$. Due to ˜3, we have that $b_{1}\in S$ and $c_{2}\in S$.
Let $d_{1}\in V(P_{1})$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Let $d_{2}\in V(P_{2})$ be the neighbor of $c_{2}$ distinct from $b_{2}$. For each $i\in\{1,2\}$, observe that $d_{i}\in T$, otherwise $d_{i}$ will be adjacent to $a_{1}$ or $b_{2}$ contradicting the fact that $C$ is an induced cycle. Now apply ˜1 on each of the edges in $\{aa_{1},a_{2}b_{2},b_{1}a_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},b_{2}^{\prime},a^{\prime\prime}_{1},b^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}a^{\prime}_{2},b_{1}a^{\prime\prime}_{1},c_{2}b^{\prime\prime}_{2}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{b^{\prime}_{2},b^{\prime\prime}_{2},b_{2}\}\subseteq B$. Similarly, apply ˜1 on edges in $\{b_{1}d_{1},c_{2}d_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},c^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}d_{1},d_{2}c^{\prime}_{2},b_{1},b^{\prime}_{1},c_{2}c^{\prime}_{2}\}\subseteq E(G)$.
@@
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
@@
Assume $b_{2}\in T$. See Figure˜3(c) for notations. Let $z$ be the vertex in $U$ which is also adjacent to $a_{2}$ in $H$. Due to ˜5, we have that $z$ is distinct from $a_{1}$. Moreover, due to ˜3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,z\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},z\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exists a $(x,a_{1})$-path $P\subseteq C$ that does not contain $a$ or $a_{2}$. Moreover, $a_{1}b_{1}\in E(P)$ and $z\notin V(P)$. Due to ˜3, we have that $b_{1}\in S$. Let $c\in V(P)$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Since $H$ is triangle-free, $c$ is not adjacent to $a_{1}$. Due to ˜5, we have $c\neq z$ and moreover, $c$ is not adjacent to $z$ in $H$. Hence, $c\notin U,c\notin S$. Therefore, $c$ must lie in $T$.
Now apply ˜1 on each of the edges in $\{aa_{1},a_{2}z,b_{1}a_{1}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},z^{\prime},a^{\prime\prime}_{1}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}z^{\prime},b_{1}a^{\prime\prime}_{1}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{z,z^{\prime}\}\subseteq B$. Similarly, apply ˜1 on edges in $\{b_{1}c,a_{2}b_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},a^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c,a^{\prime}_{2}b_{2},b_{1}b^{\prime}_{1},a_{2}a^{\prime}_{2}\}\subseteq E(G)$.
@@
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
Claims 16.1, 16.2, and 16.3 imply that $\{a,a_{2},b_{2}\}\subseteq S$. Since $\{aa_{2},a_{2}b_{2}\}\in E(H)$, we have a contradiction due to ˜5. ∎